17. A rectangular solid whose dimensions are 4cm, 8cm, and 12cm is inscribed in a sphere. What is the diagonal of the rectangular solid?

a) 7.49 cm
b) 8.94 cm
c) 14.42 cm
d) 14.97 cm
1
Expert's answer
2020-12-17T06:47:33-0500

Let "a=" the length of a rectangular solid, "b=" its width, a "c=" its height.

The the diagonal "d" of a rectangular solid will be


"d=\sqrt{a^2 +b^2+c^2}""d=\sqrt{4^2+8^2+12^2}=4\sqrt{14}(cm)\approx14.97(cm)"


d) 14.97 cm